OpenStudy (anonymous):

The number of Math homework questions is seven more than one−half the number of History homework questions. There are 15 Math homework questions. Which of the following equations would you use to find the number of History homework questions? Note: h represents the number of questions on your History homework.

OpenStudy (debbieg):

Don't worry about the answer choices, lets just figure out what the equation should be. "seven more than one−half the number of History homework questions" How do you write "one-half of {something}" as a mathematical expression? You can use "h" for the "something".

OpenStudy (anonymous):

I always write : 7+1/2h but apparently im wrong ..

terenzreignz (terenzreignz):

Probably because you need an *equation* and not just an expression.

OpenStudy (anonymous):

its multiple choice . im not sure how to even begin .

terenzreignz (terenzreignz):

You really shouldn't need the choices, but, out of interest, what are they?

OpenStudy (anonymous):

15 over 2+7=h 15+7 all over 2=h h+7 all over 2=15 1over2h + 7 = 15

OpenStudy (debbieg):

@terenzreignz , I was going to get to the equation but it doesn't hurt to do it one piece at a time. Sorry, I was in another window and didn't see the updates. You said 7+1/2h for "7 more than 1/2 of h" - that will do fine.

terenzreignz (terenzreignz):

I was just answering a concern about something being wrong all the time...

OpenStudy (anonymous):

its difficult for me to understand how to even start with this problem , but i guess 7+1/2h is a good start ,

OpenStudy (debbieg):

Now you know that: The number of Math homework questions \(\huge \text {is}\) seven more than one−half the number of History homework questions. so that quantity you just cooked up IS the number of math questions. You also know: "There are 15 Math homework questions." So how can you turn your expression into an equation? Set {what} equal to {what}?

OpenStudy (anonymous):

this is probably why i havent passed algebra 1 .. umm .. math is 15 , so .. ugh .. i dont know . i give up ..

OpenStudy (debbieg):

You are psyching yourself out! You're so close. Take a deep breath and re-read what I wrote if you need to. :)

OpenStudy (anonymous):

im guessing 1/2h+7=15 ?

OpenStudy (debbieg):

That's no guess, you are reasoning it through and that is exactly right. :)

OpenStudy (debbieg):

Your expression 7+1/2h represents the number of math problems. and you are TOLD that number IS 15. So the expression is = 15. There's your equation!

OpenStudy (debbieg):

Keeping in mind, of course, that \[7+\frac{ 1 }{ 2 }h=\frac{ 1 }{ 2 }h+7\] It's a sum, so the order of the terms doesn't matter.

OpenStudy (anonymous):

im still not undrstanding how you got that . :(

OpenStudy (debbieg):

How I got what? You got it all. :) but I'll go through it again step by step:

OpenStudy (anonymous):

so will my answer be 1/2h+7=15 ?

OpenStudy (debbieg):

The number of Math homework questions is seven more than one−half the number of History homework questions. There are 15 Math homework questions. Break it apart. "one−half the number of History homework questions" How do you write, mathematically, an expression for 1/2 of something? You multiply by 1/2 (or divide by 2, that's the same thing, but we'll stick with multiplication here). So "one-half of {something}" means \(\dfrac{1}{2}(something)\) So "one-half the number of History problems" means \(\dfrac{1}{2}h\) since you were told in the problem to use h to represent "the number of history problems". Good so far?

OpenStudy (debbieg):

Now you want: "seven more than one−half the number of History homework questions" How do you write, mathematically, an expression for seven more than {something}? You right: 7+{something}. We already know the {something} here, is the part we already came up with: So "seven more than one−half the number of History homework questions" means \(7+\dfrac{1}{2}h\)

OpenStudy (debbieg):

The last piece now. "The number of Math homework questions \(\Huge is\) seven more than one−half the number of History homework questions." So now we have that {#math questions}=seven more than one−half the number of History homework questions, that is, \( \text{#math questions}=7+\dfrac{1}{2}h\)

OpenStudy (anonymous):

got it so far .

OpenStudy (debbieg):

But the next thing they tell us is " There are 15 Math homework questions." so \(\text{#math questions}=15=7+\dfrac{1}{2}h\) giving you the final equation: \(15=7+\dfrac{1}{2}h\)

OpenStudy (debbieg):

Just break the language down one step at a time. :) re-read the sentence as many times as necessary. Figure out what the variable is and what quantity it stands in for in the problem. Then go from there.

OpenStudy (anonymous):

so i was right !!

OpenStudy (debbieg):

Yes, I told you that already! lol

OpenStudy (anonymous):

im sorry , im a junior STILL taking algebra 1 , :c its hard for me to understand sometimes .

OpenStudy (debbieg):

Well, at least you are working on TRYING to understand. That's the most important thing. If you don't work through the things you DON'T understand NOW, you will only struggle more as you progress through your math education. TRUST me.

OpenStudy (anonymous):

Oh i belive it , thank you for the assistance ! it really means a lot to me .

OpenStudy (debbieg):

As you work on it and PRACTICE it more, many of these skills and concepts will become more intuitive. And yes, I said PRACTICE because doing math is a SKILL just like playing basketball or a musical instrument. The more you practice the skill, the better you will become. So if you are struggling with a concept, seek out MORE PROBLEMS to do related to that concept! :) and you are welcome, happy to help. :)